# Simulation of the magnetic Ginzburg-Landau equation via vortex tracking

### v1

@@ Abstract.
This paper deals with the numerical simulation of the 2D magnetic time-dependent Ginzburg-Landau (TDGL) equations in the regime of small but finite (inverse) Ginzburg-Landau parameter $\epsilon$ and constant (order $1$ in $\epsilon$) applied magnetic field. In this regime, a well-known feature of the TDGL equation is the appearance of quantized vortices with core size of order $\epsilon$. Moreover, in the singular limit $\epsilon\searrow 0$, these vortices evolve according to an explicit ODE system. In this work, we first introduce a new numerical method for the numerical integration of this limiting ODE system, which requires to solve a linear second order PDE at each time step. We also provide a rigorous theoretical justification for this method that applies to a general class of 2D domains. We then develop and analyze a numerical strategy based on the finite-dimensional ODE system to efficiently simulate the infinite-dimensional TDGL equations in the presence of a constant external magnetic field and for small, but finite, $\epsilon$. This method allows us to avoid resolving the $\epsilon$-scale when solving the TDGL equations, where small values of $\epsilon$ typically require very fine meshes and time steps. We provide numerical examples on a few test cases and justify the accuracy of the method with numerical investigations.
@@ 1.3. Outline of the paper
The paper is structured as follows. In the next section, we introduce the necessary notations for precisely stating the TDGL equations and highlight some of its basic properties. We then recall some rigorous results on the asymptotic behavior of solutions to the TDGL equations in the limit where the inverse Ginzburg-Landau parameter $\epsilon$ vanishes and the applied magnetic field is kept constant. In Section 3, we describe the numerical scheme for simulating the asymptotic dynamics of vortices. In addition, we provide a theoretical justification for this numerical scheme, present numerical simulations on the vortex trajectories, and numerically analyze the convergence rates with respect to the discretization parameters. In Section 4, we present the numerical scheme to efficiently approximate solutions of the TDGL equation in the regime of small but finite $\epsilon$. We then illustrate our method with some numerical experiments. Finally, we compare the results with reference solutions computed via finite elements to numerically justify the accuracy of our method.
@@ 2. The time-dependent magnetic Ginzburg-Landau equation
At first sight, it may seem that the electric potential $\Phi$ is an additional variable of the problem. However, note that by taking the divergence of (2.3), the electric potential is, up to a constant, uniquely determined as the solution of the following Poisson equation with Neumann boundary conditions:
@@
for a (sufficiently regular) function $\chi:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$. Nevertheless, note that all of the physical quantities of interest for the model, namely the density $|u|^{2}$, the super-current $j_{A}(u)=(\imath u,\nabla_{A}u)$, the induced electric field $E=\partial_{t}A+\nabla\Phi$, and the induced magnetic field $h=\mathrm{curl}\,A$, are invariant under this gauge transformation. In particular, we are free to choose the most convenient gauge for calculations. Following [Spi03, KS11], we shall work with the Coulomb gauge here, i.e., we impose the additional conditions:
@@ 2.1. Asymptotic regime and limiting dynamics
$W_{\Omega}(a,d;h_{\rm ex})=\lim_{\rho\searrow 0}\inf_{\begin{subarray}{c}u\in\mathcal{H}^{1}_{\rho}(a)\\ A\in\mathrm{H}^{1}_{\rm div}(\Omega)\end{subarray}}\left\{\frac{1}{2}\int_{\Omega_{\rho}(a)}|\nabla_{A}u|^{2}+|\mathrm{curl}\,A-h_{\rm ex}|^{2}\mathrm{d}x\right\}-\pi\sum_{j=1}^{n}d_{j}^{2}\log\frac{1}{\rho}$
@@
which is well-defined for any non-vanishing $u\in\mathrm{H}^{\frac{1}{2}}(\partial B_{\rho}(a_{j});\mathbb{C})$ (see, e.g., [BM21]). By the results in [Spi03, Appendix A.1] (see also [LD97, Section 4.1]), the renormalized energy can be written more explicitly [^2] Here we adopt the sign convention for $\Xi$ in [KS11]. Note that this is the opposite sign convention from [Spi03, LD97]; in particular, the terms depending on $\Xi$ in the renormalized energy appear with the opposite sign there. as
@@
Strictly speaking, only the case $\alpha_{0},\beta_{0}>0$ are treated in [KS11]. The case $\alpha_{0}>0$ and $\beta_{0}=0$ is addressed in [Spi02], and the case $\alpha_{0}=0$ and $\beta_{0}>0$ can be found in [Spi03].
@@ 3.1. The modified biharmonic equation
$\int_{\Omega}\Delta\Xi_{h}(\Delta-1)\psi=-\int_{\partial\Omega}\partial_{\nu}\psi(x)\left(h_{\rm ex}+\Delta\Xi_{p}(x)\right)\mathcal{H}^{1}(\mathrm{d}x),\quad\mbox{for any $\psi\in\mathrm{H}^{1}_{0}(\Omega)\cap\mathrm{H}^{2}(\Omega)$,}$ (3.9)
where $\partial_{\nu}\psi$ denotes the Neumann trace of $\psi$ along $\partial\Omega$. The existence and uniqueness of solutions of (3.9) then holds in any bounded Lipschitz domain $\Omega$.
@@
Let $F\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.8) and $R\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.3). Since the boundary functions are smooth, it follows from standard elliptic regularity that $F,R\in\mathrm{H}^{2}(\Omega)$. Thus, let us define $\widetilde{\Xi}_{h}=F-R+h_{\rm ex}\in\mathrm{H}^{2}(\Omega)$. Then in the distributional sense we have
@@ 3.3. Discretization of the modified biharmonic equation
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded simply connected domain with Lipschitz boundary. Then the families $\{P_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ respectively $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ centered at any point $x_{0}\in\Omega$ are linearly independent and their spans are dense in $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$.
@@ 3.4. Numerical simulation of the limiting ODE
We are now capable of simulating some trajectories. In the limiting ODE (3.1), we fix $\alpha_{0}=0$, $\beta_{0}=1$ so that we are in the case of the Schrödinger flow. This ODE is solved with a 4th-order Runge–Kutta method (RK4, see e.g. [HWN93, Chapter II]) with time step $\delta t=10^{-3}$ up to some time $T=1.0$. The gradient of the renormalized energy is computed with the simplified formulation (3.13), where $F$ solves (3.8a)-(3.8b) and is approximated with $m=64$ Fourier modes on the boundary $\partial\Omega$. We consider 4 different initial conditions, each of them being used twice with $h_{\rm ex}=3$ and $h_{\rm ex}=-2$. Note that all the trajectories were obtained in a few seconds on a personal laptop: a significant improvement when compared to the computational cost of the complete PDE (2.16) for small $\epsilon$.
Such trajectories can already be analyzed from a physics perspective: in the case of vortices with positive degree, adding a positive magnetic field $h_{\rm ex}>0$ brings an additional counterclockwise rotation speed to the vortices (as can be seen from Case 1 without magnetic field and Case 1 with $h_{\rm ex}=3$ where the vortices span a shorter distance within the same time), while adding a negative magnetic field $h_{\rm ex}<0$ brings an additional clockwise rotation speed (as can be seen from Case 2 without magnetic field and Case 2 with $h_{\rm ex}=-2$). In the next section, we present how to recover an approximate solution to the PDE from the trajectories obtained in the singular limit.
@@ 4.1. Complements on well-prepared states
together with the boundary conditions
$f_{\epsilon,r_{0}}(0)=0\quad\text{and}\quad f_{\epsilon,r_{0}}(r_{0})=1.$ (4.2)
For finite $r_{0}$, this equation can be solved numerically with very high precision (e.g. using a nonlinear finite element solver), even for $\epsilon$ as small as $10^{-3}$, see Figure 3. Finally, note that, by a scaling argument, $f_{\epsilon,r_{0}}$ only depends on the ratio $r_{0}/\epsilon$, so that we may write only $f_{\epsilon}$, see [KS11, Section 5] for more details.
@@ 4.3. Vortex positions as initial conditions
Finally, the solution to Laplace’s equation (4.5) in the smoothing process at time $t>0$ is computed using harmonic polynomials of degree $m=128$. Here, the initial vortex positions $a^{0}$ are provided as input data and the initial phase $\varphi$ is computed in order to match the Neumann boundary conditions (see (4.3)). We present the numerical results obtained by the application of our numerical method to cases 1, 2, 3 and 4 from Section 3.4, for $\epsilon=0.03$ and display, for each of them, the modulus and the phase of $u_{\epsilon}^{*}$ as well as the magnetic field at different times $t$. Note that the boundary conditions of the magnetic field $h_{*}$ do match the constant value of $h_{\rm ex}$ and that, independently of the sign of $h_{\rm ex}$, vortices with a positive (resp. negative) degree induce a positive (resp. negative) magnetic field.
@@ 4.4.1. Computational framework
In order to compare the numerical method presented in Section 4.2 to the exact solution of the TDGL equation (2.16), we compute a reference solution $u_{\epsilon}(t)$ with the finite element solver Gridap.jl [BV20] and meshing tool GMSH [GR09]. The magnetic TDGL equation (2.16) is solved for initial conditions given by the vortex configuration from Case 1 and the smoothing procedure described in Section 4.1. We use a forward Euler time discretization scheme and follow a numerical scheme adapted from [GX23]. The mesh size is of order $\delta x=5\cdot 10^{-2}$ with local a priori refined mesh on the vortex trajectory with mesh size $\delta x=3.5\cdot 10^{-3}$, see Figure 13. The time step is $\delta t=10^{-5}$. The finite element solver is then used to obtain various reference solutions $u_{\epsilon}$ for $\epsilon\in\{0.07,0.1,0.13\}$, up to $T=1$. We also compute a reference solution for $\epsilon=0.05$, with a finer refinement. Note that the simulation with such a mesh can take several days for small values of $\epsilon$: for instance, in our case, the simulations for $\epsilon=0.05$ took about ten days to run with a non-optimized code on a small sized cluster.
@@ 5. Conclusion
In this work, we first presented a numerical method to simulate the limiting ODE dynamics of point-like magnetic vortices arising from the TDGL equations with constant magnetic fields. From these trajectories, we proposed a numerical method to recover approximate solutions of the TDGL equations in the regime of small but finite $\epsilon$ under a constant external magnetic field. This method avoids the use of very fine meshes to properly capture the vortices core, thereby allowing for significantly faster and accurate simulations of the physical quantities of interest such as the super current, the magnetic field, and the density of Cooper pairs. Remarkably, the method becomes more accurate the smaller $\epsilon$ is. However, the accuracy is only guaranteed as long as the vortex pairs are sufficiently separated. In particular, this prevents accurate simulations of some physical phenomena such as radiation waves emitted by vortex-antivortex annihilation, which is left for future work.

### v2

@@ Abstract.
This paper deals with the numerical simulation of the 2D magnetic time-dependent Ginzburg-Landau (TDGL) equations in the regime of small but finite (inverse) Ginzburg-Landau parameter $\epsilon$ and constant (order $1$ in $\epsilon$) applied magnetic field. In this regime, a well-known feature of the TDGL equation is the appearance of quantized vortices with core size of order $\epsilon$. Moreover, in the singular limit $\epsilon\searrow 0$, these vortices evolve according to an explicit ODE system. In this work, we first introduce a new numerical method for the numerical integration of this limiting ODE system, which requires to solve a linear second order PDE at each time step. We also provide a rigorous theoretical justification for this method that applies to a general class of 2D domains. We then develop and analyze a numerical strategy based on the finite-dimensional ODE system to efficiently simulate the infinite-dimensional TDGL equations in the presence of a constant external magnetic field and for small, but finite, $\epsilon$. This method allows us to avoid resolving the $\epsilon$-scale when solving the TDGL equations, where small values of $\epsilon$ typically require very fine meshes and time steps. We provide numerical examples on a few test cases and justify the accuracy of the method with numerical investigations. We end the paper showing that, in the mixed flow case, the limiting ODE system is able to capture the crystallization process in which, for large times, the vortices arrange into a stable pattern.
@@ 1.3. Outline of the paper
The paper is structured as follows. In the next section, we introduce the necessary notations for precisely stating the TDGL equations and highlight some of its basic properties. We then recall some rigorous results on the asymptotic behavior of solutions to the TDGL equations in the limit where the inverse Ginzburg-Landau parameter $\epsilon$ vanishes and the applied magnetic field is kept constant. In Section 3, we describe the numerical scheme for simulating the asymptotic dynamics of vortices. In addition, we provide a theoretical justification for this numerical scheme, present numerical simulations on the vortex trajectories, and numerically analyze the convergence rates with respect to the discretization parameters. In Section 4, we present the numerical scheme to efficiently approximate solutions of the TDGL equation in the regime of small but finite $\epsilon$. We then illustrate our method with some numerical experiments, and we compare the results with reference solutions computed via finite elements to numerically justify the accuracy of our method. Finally, in Section 5 we show that the limiting ODE system under a mixed-flow – i.e. adding dissipation – is capable of producing vortices that arrange after some time into a stable lattice pattern, provided that the external magnetic field is large enough.
@@ 2. The time-dependent magnetic Ginzburg-Landau equation
At first sight, it may seem that the electric potential $\Phi$ is an additional variable of the problem. However, note that by taking the divergence of (2.3), the electric potential is, up to a constant, uniquely determined as the solution of the following Poisson equation with Neumann boundary conditions, where $j_{A}(u)=(\imath u,\nabla_{A}u)$ is the super-current:
@@
for a (sufficiently regular) function $\chi:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$. Nevertheless, note that all of the physical quantities of interest for the model, namely the density $|u|^{2}$, the super-current $j_{A}(u)$, the induced electric field $E=\partial_{t}A+\nabla\Phi$, and the induced magnetic field $h=\mathrm{curl}\,A$, are invariant under this gauge transformation. In particular, we are free to choose the most convenient gauge for calculations. Following [Spi03, KS11], we shall work with the Coulomb gauge here, i.e., we impose the additional conditions:
@@ 2.1. Asymptotic regime and limiting dynamics
$W_{\Omega}(a,d;h_{\rm ex})=\lim_{\rho\searrow 0}\inf_{\begin{subarray}{c}u\in\mathcal{H}^{1}_{\rho}(a)\\ A\in\mathrm{H}^{1}_{\rm div}(\Omega)\end{subarray}}\left\{\frac{1}{2}\int_{\Omega_{\rho}(a)}|\nabla_{A}u|^{2}+\int_{\Omega}|\mathrm{curl}\,A-h_{\rm ex}|^{2}\mathrm{d}x\right\}-\pi\sum_{j=1}^{n}d_{j}^{2}\log\frac{1}{\rho}$ (2.7)
@@
which is well-defined for any non-vanishing $u\in\mathrm{H}^{\frac{1}{2}}(\partial B_{\rho}(a_{j});\mathbb{C})$ (see, e.g., [BM21]). By the results in [Spi03, Appendix A.1] (see also [LD97, Section 4.1]), the renormalized energy can be written more explicitly [^2] Here we adopt the sign convention for $\Xi$ in [KS11]. Note that this is the opposite sign convention from [Spi03, LD97]; in particular, the terms depending on $\Xi$ in the renormalized energy appear with the opposite sign there. as (see Appendix A for more details)
@@
Strictly speaking, only the case $\alpha_{0},\beta_{0}>0$ (mixed-flow) is treated in [KS11]. The case $\alpha_{0}>0$ and $\beta_{0}=0$ (heat-flow) is addressed in [Spi02], and the case $\alpha_{0}=0$ and $\beta_{0}>0$ (Schrödinger-flow) can be found in [Spi03].
@@ 3.1. The modified biharmonic equation
$\int_{\Omega}\Delta\Xi_{h}(\Delta-1)\psi=-\int_{\partial\Omega}\partial_{\nu}\psi(x)\left(h_{\rm ex}+\Delta\Xi_{p}(x)\right)\mathscr{H}^{1}(\mathrm{d}x),\quad\mbox{for any $\psi\in\mathrm{H}^{1}_{0}(\Omega)\cap\mathrm{H}^{2}(\Omega)$,}$ (3.9)
where $\partial_{\nu}\psi$ denotes the Neumann trace of $\psi$ along $\partial\Omega$. The existence and uniqueness of solutions of (3.9) then holds in any bounded Lipschitz domain $\Omega$, see Appendix B for more details.
@@
Let $F\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.8) and $R\in\mathrm{H}^{1}(\Omega)$ be a weak solution of (3.3). Since the boundary functions are smooth on a neighbourhood of the boundary, it follows from standard elliptic regularity that $F,R\in\mathrm{H}^{2}(\Omega)$. Thus, let us define $\widetilde{\Xi}_{h}=F-R+h_{\rm ex}\in\mathrm{H}^{2}(\Omega)$. Then in the distributional sense we have
@@ 3.3. Discretization of the modified biharmonic equation
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded simply connected domain with Lipschitz boundary. Then the families $\{P_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$ (resp. $\{Q_{j}\rvert_{\partial\Omega}\}_{j\in\mathbb{Z}}$) centered at any point $x_{0}\in\Omega$ are linearly independent and their spans are dense in $\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$.
@@ 3.4. Numerical simulation of the limiting ODE
We are now capable of simulating some trajectories. In the limiting ODE (3.1), we fix $\alpha_{0}=0$, $\beta_{0}=1$ so that we are in the case of the Schrödinger flow. This ODE is solved with a 4th-order Runge–Kutta method (RK4, see e.g. [HWN93, Chapter II]) with time step $\delta t=10^{-3}$ up to some time $T$. The gradient of the renormalized energy is computed with the simplified formulation (3.13), where $F$ solves (3.8a)-(3.8b) and is approximated with $m=64$ Fourier modes on the boundary $\partial\Omega$. We consider 4 different initial conditions, each of them being used twice with $h_{\rm ex}=3$ and $h_{\rm ex}=-2$. Note that all the trajectories were obtained in a few seconds on a personal laptop: a significant improvement when compared to the computational cost of the complete PDE (2.17) for small $\epsilon$.
Such trajectories can already be analyzed from a physics perspective: in the case of vortices with positive degree, adding a positive magnetic field $h_{\rm ex}>0$ brings an additional counterclockwise rotation speed to the vortices (as can be seen from Case 1 without magnetic field and Case 1 with $h_{\rm ex}=3$ where the vortices span a shorter distance within the same time), while adding a negative magnetic field $h_{\rm ex}<0$ brings an additional clockwise rotation speed (as can be seen from Case 2 without magnetic field and Case 2 with $h_{\rm ex}=-2$).
@@
In the next section, we present how to recover an approximate solution to the PDE from the trajectories obtained in the singular limit.
@@ 4.1. Complements on well-prepared states
together with the boundary conditions $f_{\epsilon,r_{0}}(0)=0$ and $f_{\epsilon,r_{0}}(r_{0})=1$. For finite $r_{0}$, this equation can be solved numerically with very high precision (e.g. using a nonlinear finite element solver), even for $\epsilon$ as small as $10^{-3}$, see Figure 3. Finally, note that, by a scaling argument, $f_{\epsilon,r_{0}}$ only depends on the ratio $r_{0}/\epsilon$, so that we may write only $f_{\epsilon}$, see [KS11, Section 5] for more details.
@@ 4.3. Vortex positions as initial conditions
Finally, the solution to Laplace’s equation (4.4) in the smoothing process at time $t>0$ is computed using harmonic polynomials of degree $m=128$. Here, the initial vortex positions $a^{0}$ are provided as input data and the initial phase $\varphi$ is computed in order to match the Neumann boundary conditions (see (4.2)). We present the numerical results obtained by the application of our numerical method to cases 1, 2, 3 and 4 from Section 3.4, for $\epsilon=0.02$ and display, for each of them, the modulus and the phase of $u_{\epsilon}^{*}$ as well as the magnetic field at different times $t$. Note that the boundary conditions of the magnetic field $h_{*}$ do match the constant value of $h_{\rm ex}$ and that, independently of the sign of $h_{\rm ex}$, vortices with a positive (resp. negative) degree induce a positive (resp. negative) magnetic field.
@@ 4.4.1. Computational framework
In order to compare the numerical method presented in Section 4.2 to the exact solution of the TDGL equation (2.17), we compute a reference solution $u_{\epsilon}(t)$ with the finite element solver Gridap.jl [BV20] and meshing tool GMSH [GR09]. The magnetic TDGL equation (2.17) is solved for initial conditions given by the vortex configuration from Case 1 and the smoothing procedure described in Section 4.1. We use a forward Euler time discretization scheme and follow a numerical scheme adapted from [GX23]. The mesh size is of order $\delta x=5\cdot 10^{-2}$ with local a priori refined mesh on the vortex trajectory with mesh size $\delta x=3.5\cdot 10^{-3}$, see Figure 13. The time step is $\delta t=10^{-5}$. The finite element solver is then used to obtain various reference solutions $u_{\epsilon}$ for $\epsilon\in\{0.07,0.1,0.13\}$, up to $T=1$. We also compute a reference solution for $\epsilon=0.05$, with a finer refinement. For this finer mesh, the total number of degrees of freedom for $u_{\epsilon}$, $A_{\epsilon}$ and $\varphi_{\epsilon}$ together is 1.422.686. Note that the simulation with such a mesh can take several days for small values of $\epsilon$: for instance, in our case, the simulations for $\epsilon=0.05$ took about ten days to run with a non-optimized code on a small sized cluster.
@@ 4.4.2. Numerical illustration of the singular limit $\epsilon\searrow 0$
5. The mixed-flow case
To conclude, we end this paper by simulating the ODE system (3.1) in the mixed flow case, that is the dissipative coefficient $\alpha_{0}$ is nonzero. In this case, it is expected that, for a large enough external magnetic field $h_{\rm ex}$, the dissipation forces the vortices to arrange, for large times, into a lattice pattern [SS07]. In Figure 17 we plot, for four different cases, the limiting pattern observed at time $T=10$. The discretization parameters of the ODE solver described in Section 3.4 are $\delta t=10^{-3}$ with $m=64$ Fourier modes on the boundary. The initial conditions are given by random positions inside the disc of $n\in\{10,30,50,80\}$ vortices with same degree +1+1. As expected, the vortices align according to some pattern for large times. Note also that the more numerous the vortices, the larger the magnetic field $h_{\rm ex}$ has to be in order to confine the vortices inside the domain $\Omega$: for instance, with $n=80$ vortices, if $h_{\rm ex}=100$, we observe numerically that the vortices escape from $\Omega$ in finite time.
Moreover, we present in Figures 1821 the approximate solutions to the full PDE in the mixed-flow case for $\epsilon=0.01$ at different time $t$ obtained with the method introduced in Section 4.2, showing how the vortices are converging towards the lattice pattern.
Case $n$ $h_{\rm ex}$ $\alpha_{0}$ $\beta_{0}$ vortex degree
Case 1 10 50 1 1 +1+1 for all
Case 2 30 100 1 5 +1+1 for all
Case 3 50 150 1 10 +1+1 for all
Case 4 80 250 1 10 +1+1 for all
Table 5.1. Initial configurations for simulation of the ODE system (3.1) in the mixed-flow case.
Figure 17. Asymptotic pattern of point vortices under the mixed-flow ODE (3.1). From left to right, top to bottom: case 1, case 2, case 3 and case 4 from Table 5.1.
Figure 18. Squared modulus and phase of $u_{\epsilon}^{*}(t)$ at various times $t$ under a mixed-flow (Case 1).
Figure 19. Squared modulus and phase of $u_{\epsilon}^{*}(t)$ at various times $t$ under a mixed-flow (Case 2).
Figure 20. Squared modulus and phase of $u_{\epsilon}^{*}(t)$ at various times $t$ under a mixed-flow (Case 3).
Figure 21. Squared modulus and phase of $u_{\epsilon}^{*}(t)$ at various times $t$ under a mixed-flow (Case 4).
@@ 6. Conclusion
In this work, we first presented a numerical method to simulate the limiting ODE dynamics of point-like magnetic vortices arising from the TDGL equations with constant magnetic fields. From these trajectories, we proposed a numerical method to recover approximate solutions of the TDGL equations in the regime of small but finite $\epsilon$ under a constant external magnetic field and applied it for both the Schrödinger-flow and mixed-flow cases. This method avoids the use of very fine meshes to properly capture the vortices core, thereby allowing for significantly faster and accurate simulations of the physical quantities of interest such as the super current, the magnetic field, and the density of Cooper pairs. Remarkably, the method becomes more accurate the smaller $\epsilon$ is. However, the accuracy is only guaranteed as long as the vortex pairs are sufficiently separated. In particular, this prevents accurate simulations of some physical phenomena such as radiation waves emitted by vortex-antivortex annihilation, which is left for future work.
Data availability
All the codes used to generate the plots and run the simulations from this paper are available at
https://doi.org/10.18419/DARUS-5604